OpenStudy (anonymous):

Let q be the number of quarters, d be the number of dimes, and n be the number of nickels.

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

25q + 10d + 5n + 60 = 187 q + d + n + 60 = 187 25q + 10d + 5n + 60 = 1.87 0.25q + 0.10d + 0.5n + 0.60 = 187